A company estimates that 7% of their products will fail after the original warranty period but within 2 years of the purchase, with a replacement cost of $250.


If they want to offer a 2 year extended warranty, what price should they charge so that they'll break even (in other words, so the expected value will be 0)


1
Expert's answer
2021-10-25T18:09:28-0400

A company estimate is 7%

A replacement cost of $250.

"Charge = 250 \times \frac{7}{100} \\

= 250 \times 0.07 \\

= \$17.5"